Two systems have the following equations of state and are contained in a closed cylinder, separated by a fixed, adiabatic and impermeable piston. N1=2, N2=1.5
. The initial temperatures are T1 = 175 K and T2 = 400 K. The total volume is 0.025 m^3
. The piston is allowed to move and heat transfer is allowed across the piston. Determine the final temperature of the system (in Kelvin).
1/T1=3/2(R*N1/U1) , P1/T1=R*N1/V1
1/T2=5/2(R*N2/U2) , P2/T2=R*N2/V2
1
Expert's answer
2020-06-29T08:29:27-0400

Let "V_1 ,V_2" are the volume of two system

Now,

"V_1+V_2= V= 0.025 m^3"

As both system are in closed container in adiabatic condition

"T_1=175 K,T_2=400 K"

we know that, PV=nRT

"\frac{nRT_1}{P_1}+\frac{nRT_2}{P_2}= \frac{nRT}{P}"


"\frac{2R\times 175}{P}+\frac{1.5R\times 400}{P}= \frac{3.5RT}{P}"

T= 270.42 K, This is the final temeprature